OpenStudy (anonymous):

What does it mean to refer to a 20-in TV or 25-in TV set? Such units refer to diagonal of the screen. A 15-in TV set also has the width of 12 in. What is the height?

jimthompson5910 (jim_thompson5910):

a^2+b^2 = c^2 ... the pythagorean theorem 12^2+b^2 = 15^2 ... a = 12 (the width) and c = 15 (the diagonal) From here, we solve for b 144 + b^2 = 225 b^2 = 225 - 144 b^2 = 81 b = sqrt(81) ... Note: all lengths are positive, so we're dealing with the positive square root b = 9 So the height is 9 inches